You suspect that chloral hydrate has been used to “knock out” one of your colleagues due to its characteristic odour. A samplPlease show all work! Thanks. The right answer should be C2H3O4Cl3.
Answer #1

from the the given amount of information we have to find C, H, 0 and ce. Moles o CO₂ = Mass of co2 - Molar mass = 0.13219 = 0О. ISBS + с). Mass of ce = 0.0045 melx 35.45 gmaly. --1 Sss a + Момо ҳам — (мом о кари) - ( мам в сян = (0-2 843 4) - (oo36 aEmberi cal formula mass = (12+ 105 X1 +1.5x39.45 + 1758 16) glmel = 94.675 g limel. Molecular formula = nx Emperical formula

Right answer is C4H6Cl6O7 not C2H3Cl3O4